OpenStudy (dragon123mama):

evaluate the logarithm log 6 (1/36)

OpenStudy (dragon123mama):

@mayankdevnani

OpenStudy (dragon123mama):

\[\log{6} \frac{ 1 }{ 36 }\]

TheSmartOne (thesmartone):

\(\LARGE log_\color{red}{a}\color{blue}{ b} = \color{green}{c}\) \(\LARGE \color{red}{a}^{\color{green}{c}} = \color{blue}{b}\)

OpenStudy (dragon123mama):

oh so i just have to put it in the other form?

OpenStudy (dragon123mama):

there is no equals though so how would i do it without C?

MARC (marc_d):

\[=\log_6\frac{ 1 }{ 6^2 }\]

OpenStudy (dragon123mama):

would it just be 6=(1/36)

TheSmartOne (thesmartone):

\[\log_{6} \frac{ 1 }{ 36 }=x\] change it into exponential form and solve for x :)

OpenStudy (dragon123mama):

so 6(x)= (1/36)

TheSmartOne (thesmartone):

it should 6^x and hint: \(\Large a^{-b} = \frac{1}{a^b}\)

OpenStudy (dragon123mama):

yeah that hint doens't tell me much?

OpenStudy (dragon123mama):

so it is 6^x= (1/36)

OpenStudy (dragon123mama):

hello?

OpenStudy (dragon123mama):

okay bye then.